OpenStudy (ny,ny):

You buy chicken wire to enclose a rectangular garden. You are restricted to a width of no more than 30 feet. You would like to use at most 180 feet chicken wire. write system of linear inequalities to describe this situation.

OpenStudy (anonymous):

so then plug it in

OpenStudy (anonymous):

wait never mind

OpenStudy (ny,ny):

i meant linear inequalities. sorry.

OpenStudy (anonymous):

yeah hang on ok w< or equal to 30

OpenStudy (ny,ny):

2l + 2w < = 180 do you think?

OpenStudy (anonymous):

so the 2l+2(15)=180

OpenStudy (anonymous):

yeah you were correct

OpenStudy (anonymous):

also put up the inequality that w<= 30

OpenStudy (ny,ny):

so the equations are w < = 30 2l + 2w < = 180?

OpenStudy (anonymous):

yeah!

OpenStudy (anonymous):

if that helped make sure you let me know with best answer:)

OpenStudy (ny,ny):

do you know how to graph this?